OpenStudy (anonymous):

which contains more molecules: 1mole H2O2, 1 mole H2O, or 1 mole CO

OpenStudy (aaronq):

1 mole = 6.022*10^23 molecules (or particles, ions, cats, etc) If you were to split the molecules into individual atoms, the answer would be different.

OpenStudy (anonymous):

\[\frac { 1\quad mol\quad H_{ 2 }O_{ 2 } }{ 1 } *\frac { 6.022\ast 10^{ 23 }\quad molecules\quad H_{ 2 }O_{ 2 } }{ 1\quad mol\quad H_{ 2 }O_{ 2 } } =\quad 6.022\ast 10^{ 23 }\quad molecules\quad H_{ 2 }O_{ 2 }\] \[\frac { 1\quad mol\quad H_{ 2 }O }{ 1 } *\frac { 6.022\ast 10^{ 23 }\quad molecules\quad H_{ 2 }O }{ 1\quad mol\quad H_{ 2 }O } =\quad 6.022\ast 10^{ 23 }\quad molecules\quad H_{ 2 }O\] \[\frac { 1\quad mol\quad CO }{ 1 } *\frac { 6.022\ast 10^{ 23 }\quad molecules\quad CO }{ 1\quad mol\quad CO } =\quad 6.022\ast 10^{ 23 }\quad molecules\quad CO\]